OpenStudy (anonymous):

Solve 5 over x+7 = 3 over x+2?

mathslover (mathslover):

\[\huge{\frac{5}{x+7}=\frac{3}{x+2}}\] \[\huge{5(x+2)=3(x+7)}\] \[\huge{5x+10=3x+21}\] \[\huge{2x=11}\] \[\huge{x=\frac{11}{2}}\]